Find the integral surface for the differential equation z(xzx-yzy)=y2-x2 passing through the initial data curve (2s, s, s)


1
Expert's answer
2021-01-15T14:20:38-0500

"z(xz_x-yz_y)=y^2-x^2\\
xz_x-yz_y=\frac{y^2-x^2}{z}\\
P=x , Q=-y, R=\frac{y^2-x^2}{z}\\
\frac{dx}{dt}=P=x\\
\frac{dx}{x}=dt \text{ Integrate}\\
\ln x=t+c_1\\
x=C_1e^t\\
\text{From the initial curve, we know that } x(0)=2s\\
C_1=2s\\
x=2se^t....................(1)\\
\text{Similarly, }\\
\frac{dy}{dt}=Q=-y\\
-\ln y=t+c_2\\
y=C_2e^{-t}\\
y(0)=s\\
C_2=s\\
y=se^{-t}................(2)\\
\text{Also,}\\
\frac{dz}{dt}=R=\frac{y^2-x^2}{z}\\
zdz=(y^2-x^2)dt\\
\frac{z^2}{2}=(y^2-x^2)t+c_3\\
z(0)=s\\
c_3=\frac{s^2}{2}\\
z^2=2(y^2-x^2)t+s^2...............(3)"

We need to get the value for "s \text{ and }t"

From (1) and (2),

"\frac{x}{y}=2e^{2t}\\
t=\ln(\frac{x}{2y})^2"

Substitute this into (1)

"x=2se^{\ln (\frac{x}{2y})^2}\\
x=\frac{2s.x^2}{4y^2}\\
s=\frac{2y^2}{x}"

Put the values if s and t into (3) to get the solution.

"z^2=2(y^2-x^2)\ln (\frac{x}{2y})^2+\frac{4y^4}{x^2}\\
x^2z^2=2x^2(y^2-x^2)\ln (\frac{x}{2y})^2+4y^4"

Thus is the solution and the characteristics are

"x=\frac{2y^2}{s}"